OpenStudy (anonymous):

How do you solve this problem 5∜7- 8∛7- 2∜7- ∛7 ?

OpenStudy (anonymous):

Group the ones with the same index together:

OpenStudy (anonymous):

= 5∜7 - 2∜7 = 3 ∜7 Similarly for the others: ...

OpenStudy (lgbasallote):

\[\huge 5\sqrt[4]7 - 8\sqrt[3]7 - 2\sqrt[4]7 - \sqrt[3]7\]

OpenStudy (lgbasallote):

that right?

OpenStudy (anonymous):

Yes..

OpenStudy (lgbasallote):

so you just do what chloro said

OpenStudy (anonymous):

What do you mean by index?

OpenStudy (anonymous):

@Chlorophyll?

OpenStudy (anonymous):

whether it's square root, cube root, fourth root, and so on.

OpenStudy (anonymous):

@0mgl0la did you recognize I group ∜7 together? Why don't you mimic it by grouping the ∛7 terms together: ...